### v1

# Polarized multiwavelength emission from pulsar wind – accretion disk interaction in a transitional millisecond pulsar

Abstract
Transitional millisecond pulsars (tMSPs) bridge the evolutionary gap between accreting neutron stars in low-mass X-ray binaries and millisecond radio pulsars. These systems exhibit a unique subluminous X-ray state characterized by the presence of an accretion disk and rapid switches between high and low X-ray emission modes. The high mode features coherent millisecond pulsations spanning from the X-ray to the optical band. We present multiwavelength polarimetric observations of the tMSP PSR J10230038 aimed at conclusively identifying the physical mechanism powering its emission in the subluminous X-ray state. During the high mode, we detect polarized emission in the 2–6 keV energy range, with a polarization degree of and a polarization angle of (1) measured counterclockwise from the North celestial pole towards East. At optical wavelengths, we find a polarization degree of and a polarization angle aligned with that in the soft X-rays, suggesting a common physical mechanism operating across these bands. Remarkably, the polarized flux spectrum matches the pulsed emission spectrum from optical to X-rays. The polarization properties differ markedly from those observed in other accreting neutron stars and isolated rotation-powered pulsars and are also inconsistent with an origin in a compact jet. Our results provide direct evidence that the polarized and pulsed emissions both originate from synchrotron radiation at the shock formed where the pulsar wind interacts with the inner regions of the accretion disk.
1 Introduction
The evolution of accreting neutron stars (NSs) in low-mass X-ray binaries (LMXBs) as progenitors of millisecond radio pulsars (MSPs) involves distinct phases [1], with transitional MSPs (tMSPs) representing one of the most intriguing stages of this process due to their unique properties [23, 95]. PSR J10230038 (hereafter J1023) is the prototypical tMSP [4], offering an exceptional opportunity to study these enigmatic objects in detail. Detected in 2007 as a radio MSP with a spin period of 1.69 ms, J1023 orbits a 0.24 star every 4.75 hours [4]. In 2013, J1023 experienced dramatic changes in its emission properties, ceasing its radio pulsations and exhibiting increased emission across the electromagnetic spectrum [112, 101]. J1023 has since remained in an active state, characterized by subluminous X-ray emission with an average luminosity of erg s-1 in the 0.3–80 keV energy range [33] – over four orders of magnitude below the Eddington luminosity for a NS. This emission alternates between high-intensity (70–80% of the time) and low-intensity (20–30%) modes, with occasional brief flares reaching erg s-1 [96, 81, 5, 18]. Mode switches occur within 10–30 s, with low modes lasting from tens of seconds to minutes. During the high mode, J1023 exhibits coherent pulsed emission at the pulsar spin period at X-ray, ultraviolet (UV), and optical wavelengths, pointing to processes linked to the NS rotation [5, 98, 90, 70]. This pulsed emission disappears during the low mode. The UV, optical, and near-infrared emissions mainly originate from an accretion disk surrounding the NS and the irradiated companion star, leading to flickering and flaring [74, 97, 59, 109]. Bright radio and mm continuum emission is observed, with increased radio emission during the low mode and brief mm-band flares during switches from high to low modes [37, 19, 10].
The unique behavior of J1023 raises fundamental questions about the mechanisms powering its emission in the active state. J1023 releases rotational kinetic energy at a rate of erg s-1 [3], about six times its average X-ray luminosity. This has sparked debates on whether the emission of tMSPs like J1023 in this active state is powered by pulsar rotation, mass accretion, or a combination of both [19, 98, 122, 82]. Multiband polarization measurements provide a key diagnostic tool for investigating these processes and solving this enigma.
This paper presents the results from the first multiwavelength polarimetric campaign of J1023 and is structured as follows. Section 2 describes the observations and data processing techniques employed across the X-ray, optical, and radio bands using the Imaging X-ray Polarimetry Explorer (IXPE), the Neutron Star Interior Composition Explorer Mission (NICER), the Neil Gehrels Swift Observatory (Swift), the Very Large Telescope (VLT), and the Karl G. Jansky Very Large Array (VLA). Section 3 details the data analysis and results, encompassing the selection of emission modes and timing analysis of the IXPE data, and the polarimetric analysis across different wavelengths. We also present the spectral energy distributions (SEDs) derived from our observations. Section 4 discusses the implications of our findings on the emission mechanisms powering J1023 during its active state, evaluating scenarios such as accretion-powered processes, rotation-powered mechanisms analogous to isolated pulsars, and interactions between the pulsar wind and the accretion flow. Section 5 summarizes how our multiwavelength polarimetric observations provide critical insights into the enigmatic behavior of tMSPs like J1023.
2 The multi-wavelength dataset
We acquired multiwavelength data of J1023 using space-borne and ground-based telescopes. This section describes the observations and the procedures used for data processing and analysis (for more details, see Appendix A). All timestamps are referenced to the Solar System barycenter using the JPL DE440 ephemeris and the position: RA = 10h23m47s69, Dec = +00∘38′40.′′8 (J2000.0; [36]).
2.1 IXPE Observations
IXPE observed J1023 during two sessions: May 29–June 2 and June 3–June 14, 2024 (MJD 60459.52–60463.76; MJD 60464.88–60475.62). The total exposure time per detector unit (DU) was 675 ks. To reduce instrumental background, we applied the rejection algorithm by [40]111https://github.com/aledimarco/IXPE-background. Source photons were extracted from a circle of 60′′ radius centered on the source, and background photons from an annulus with radii of 150′′ and 300′′. Emission modes were selected as detailed in Appendix B. Due to background dominance above 6 keV, we restricted the polarimetric analysis to the 2–6 keV band. Data processing was performed using the ixpeobssim package [11]. Spectral and spectro-polarimetric analyses were conducted with HEASoft and FTOOLS using the latest calibration files.
2.2 NICER Observations
During the IXPE campaign, we conducted two NICER observations of J1023 (Table 2) to guide the selection of high and low-mode episodes in the IXPE dataset. These observations were performed during orbit night-time to avoid optical light leak due to a damaged thermal shield222https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/light-leak-overview/. We reprocessed the data using NICERDAS version 13 in HEASoft and calibration products from 2024 February 6. We calibrated and screened the data using nicerl2, and extracted 0.5–10 keV time series binned at 10 s using nicerl3-lc. The background level was estimated with the SCORPEON model333https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/scorpeon-overview/.
2.3 Swift Observations
We performed two observations with Swift simultaneously with the VLA and VLT on 2024 June 4–5 (Table 2), using the X-ray Telescope (XRT) in photon counting mode and the Ultra-Violet/Optical Telescope (UVOT) in event mode with the UVM2 filter (central wavelength 226 nm; FWHM 52.7 nm). Data were processed using calibration files from 2024 May 22 for XRT and 2024 February 1 for UVOT. For XRT, we extracted source photons using a circular aperture of 47.2′′ radius and background using an annulus with radii 94.4′′ and 188.8′′. We extracted background-subtracted 0.5–10 keV time series binned at 50 s (Fig. 9). For UVOT, we used coordinator and uvotscreen to obtain cleaned event lists, and extracted time series binned at 30 s using uvotevtlc (Fig. 9).
2.4 VLT Observations
We observed J1023 with the FOcal Reducer/low dispersion Spectrograph 2 (FORS2) on the VLT in polarimetric mode (Table 2). The observations were performed from 2024 June 4 at 23:34:37 UTC to 2024 June 5 at 00:39:23 UTC, under photometric conditions (seeing 0.3”). A total of 56 images were acquired using the filter ( band; central wavelength 655 nm; FWHM 165 nm), each with a 20 s exposure, totaling 3946 s. A Wollaston prism and rotating half-wave plate (HWP) allowed images at four angles , . Four sets of images were acquired for each angle, repeated 14 times.
The images were bias-subtracted and flat-fielded. Aperture photometry was performed using the daophot tool [114] with a 6-pixel aperture. Using Equations 1–2 of [9], we calculated the normalized Stokes parameters and . To estimate the degree and angle of linear polarization, we evaluated the parameter for each of the HWP angles, following [41] and references therein (see also [34, 9, 10]). This is linked to the the polarization degree and the angle by the formula:
The parameters and were determined by maximizing the Gaussian likelihood function to provide an initial estimate for and and refining the estimates using a Markov Chain Monte Carlo (MCMC) procedure [67] (for further details on the algorithm, see [9, 10]). The polarization angle was calibrated using observations of the polarized standard star Vela 1-95, resulting in a correction of .
2.5 VLA Observations
We obtained simultaneous VLA radio observations of J1023 with IXPE and VLT from 2024 June 4 at 23:00:00 UTC to June 5 at 02:59:20 UTC. The source underwent a bright multi-wavelength flare starting at 00:36:00 UTC on June 5. We include here only the pre-flare data, yielding about 1.2 hr on-source.
We observed in C-band (central frequency 6 GHz) with a 3-bit sampler, dividing the 4096 MHz band into 32 spectral windows with 64 2-MHz channels each. Calibration used 3C286 (flux, bandpass), J10240052 (gain), and J14072827 (leakage). J1023 was monitored in 9-min target scans bracketed by two 1-min secondary scans.
Data were processed using the CASA VLA pipeline [26], followed by manual and automated (rflag and tfcrop) radio frequency interference flagging. Cross-hand calibration followed standard VLA polarization procedures444https://casaguides.nrao.edu/index.php/CASA_Guides:Polarization_Calibration_based_on_CASA_pipeline_standard_reduction:_The_radio_galaxy_3C75-CASA6.5.4; Stokes images were created with WSClean [94]. After initial shallow deconvolution of the 10′ field of view (FOV), we created a deconvolution mask using the unmasked Stokes image, then iteratively refined imaging and performed phase-only self-calibration using QuartiCal [75]. We then subtracted background sources, allowing for narrow FOV images minimally affected by aliasing.
J1023 showed an average flux density of 240 Jy and experienced a short (8 min) mini-flare peaking at 500,Jy. We produced time-resolved images in 1-min intervals and also created four deeper polarization images by combining multiple scans to boost signal-to-noise ratio at intervals corresponding to the times before the mini-flare, the mini-flare, the times after the mini-flare, and the combined times before and after the mini-flare. For each interval, WSClean generated 16 channelized images and a single frequency-averaged Multi-Frequency Synthesis (MFS) image. Flux densities were measured using imfit in CASA, modeling the source as a point source with a synthesized beam-shaped Gaussian fit. uncertainties were derived from RMS noise estimates in nearby source-free regions using imstat, employing a circular aperture encompassing 100 synthesized beam areas. When signal-to-noise permitted, we fitted a power-law of the form to channelized flux densities to determine the spectral index .
3 Data Analysis and Results
Figure 1 shows the multiwavelength time series extracted from our dataset, along with the evolution of the optical and radio polarization properties. During the observations, we detected a single episode of low mode at X-ray and UV wavelengths lasting 5 min, matching an enhanced flux at both optical and radio wavelengths. In the following sections, we describe the multiwavelength timing and polarimetric properties of J1023 and present the SEDs of the total, pulsed, and polarized emissions.
3.1 Timing Analysis of the IXPE Dataset
Details on our timing analysis of X-ray data are provided in Appendix C. Coherent X-ray pulsations at the NS spin period in J1023 are detected only during the high mode [5, 71, 98, 70]. Hence, we used source photons collected by IXPE in the high mode in the 2–6 keV band for our analysis. We corrected photon arrival times for the pulsar’s orbital motion, using an orbital period of = 17115.5216592 s and a projected semi-major axis of asini/c = 0.343356 light-seconds [3]. We performed an epoch folding search around the expected spin frequency, sampling each spin cycle in 16 phase bins. We used a grid of values centered on the predicted value from [70], MJD. We recovered the coherent signal with an epoch folding variance of (single-trial false alarm rate ) at a spin frequency Hz and = 60461.979580(14) MJD. The uncertainties were evaluated following [80] and from the Gaussian fit to the pulse variance distribution. The timing parameters are reported in Table 3555The timing model is archived at Zenodo.. Our measurement tends to favor a scenario where the pulsar’s rotational evolution in the subluminous X-ray state remains largely unchanged compared to its behavior during the radio pulsar state (Fig. 11).
3.2 X-ray Polarimetric Properties
Details on the polarimetric analysis of the IXPE data are given in Appendix D. Within the 3–6 keV band, we detected a polarization degree of % and a polarization angle , measured counter-clockwise from the North celestial pole toward the East. This polarization degree exceeds the Minimum Detectable Polarization (MDP; [126, 45]) of 15% at 99% confidence, allowing us to reject the null hypothesis of unpolarized emission with 99.4% confidence (3.2 significance; 666https://heasarc.gsfc.nasa.gov/docs/ixpe/analysis/IXPEStats-Advice.pdf). In the 2–3 keV band, the polarization degree is consistent with zero within a 1 uncertainty (see Fig. 2). By applying the same analysis to high-mode data, we obtain values consistent with the above (see Fig. 3), though none exceeded the MDP at 99% confidence. No variations in Stokes parameters were observed across the pulsar rotational cycle or the binary orbital cycle (for details, see Appendix D.3). For low-mode and flaring-mode data, the significance was too low for a reliable analysis, yielding upper limits of % and % at 90% confidence over the 2–6 keV range (Appendix D.1).
To improve the polarimetric measurements, we performed weighted spectro-polarimetric analyses [116, 38] on both the entire dataset and the high-mode dataset (Appendix D.2). We fitted a power-law model corrected for interstellar absorption [18, 33] and convolved with a constant polarization model to the data in the 2–6, 2–3 and 3–6 keV ranges. Results are summarized in Table 1. For the average emission in the 2–6 keV range, we obtained % at . For the high-mode emission in the same range, we measured % at . Figure 4 shows the confidence contours for the measurements of and in the high mode across energy bands.
Previous analyses show that 10% of low-mode episodes in the IXPE datasets may be missed due to their short durations [5, 18, 33]. However, this does not affect the polarization degree measured during the high mode, as the potential increase, about 0.2% in the polarization degree, is within the measurement uncertainties (Appendix D.2).
Energy band | Average emission | High mode | ||||
(keV) | (%) | (∘) | (%) | (∘) | ||
2–6 | ||||||
2–3 | ||||||
3–6 |
3.3 Optical Properties
Details on the optical properties of J1023 are presented in Appendix E. J1023 was detected at an average magnitude of 16.4 (Vega) in the -band. We measured an average linear polarization degree of and a polarization angle of . During the X-ray high mode, we measured at an angle . This polarization angle is fully consistent with that derived at soft X-rays energies within the uncertainties (Fig. 4). In the X-ray low mode, we measured and , indicating a slight decrease in polarization degree during the low mode.
3.4 Radio Properties
J1023 exhibited an average flux density of 240 Jy at 6 GHz and experienced a mini-flare matching a low-mode episode in X-rays lasting 8 min, reaching a peak flux density of 500 Jy. The power-law spectral index remains nearly constant during the high-mode episodes that precede (H1) and follow (H2) the low mode, but decreases during the low mode (L): we obtain , and . Although the differences between modes do not exceed the 2 significance threshold, the consistency of this trend with previous observations [19] points to optically thick synchrotron emission in the high mode and a mix of optically thick and optically thin synchrotron emission in the low mode.
These observations also provide the first constraints on the linear polarization fraction () of the radio emission in both modes (Appendix A.5). We establish 3 upper limits of , and . By stacking high-mode data, we achieved .
3.5 Spectral Energy Distributions
To extract the SEDs shown in Fig. 5, we combined the fluxes for the emission in the high mode measured from IXPE and the VLT. Specifically, for the total SED we used the unabsorbed X-ray fluxes in the 2–3, 3–4 and 4–6 keV energy bands. For the optical data, we averaged the -band fluxes from our VLT/FORS2 dataset during high modes. Adopting the estimated hydrogen column density (), we calculated the -band absorption coefficient based on the relation from [50] and determined the -band absorption using the extinction laws from [58] (see also [57, 49, 55, 35]).
For the polarized SED, for the X-ray band we focused on the 2–3 keV and 3–6 keV energy ranges, using the polarization degrees measured during the high mode through spectro-polarimetric analyses (see Table 1). The dereddened polarized optical flux was estimated by rescaling according to the polarization degree measured in the high mode, .
For comparison, we overlaid the total and pulsed X-ray and optical fluxes obtained from a previous campaign involving the XMM-Newton and NuSTAR satellites as well as the SiFAP2 photometer mounted on the INAF Telescopio Nazionale Galileo [98]. Additionally, we included the fit of their SED of the pulsed emission using a single power law defined by , where is the flux density at frequency , is the normalization constant, and is the spectral index. Using nonlinear least squares fitting, the best-fit parameters were determined to be , and [98].
Our analysis shows that the SED of the polarized flux is consistent with the same power-law model that describes the spectrum of the pulsed signal. This provides additional strong evidence that the observed multiwavelength polarized emission shares a common origin with the pulsed emission across both optical and X-ray bands. Moreover, the near-coincidence of the polarized and pulsed fluxes throughout the spectrum indicates that the observed polarization originates almost entirely from the same emission processes responsible for the pulsations across the optical and X-ray frequencies. Theoretical predictions for the linear polarization of synchrotron radiation, assuming a spectral index (as measured by [98] and shown in Fig. 5), suggest a theoretical maximum polarization degree of 77% [107], which is consistent with our findings. Future instrumentation designed to measure the polarization of pulsed emission will be essential to test this.
4 Discussion
4.1 The Nature of the Polarized Emission
4.1.1 An Accretion-powered Mechanism?
Our observations reveal that J1023 exhibits a higher soft X-ray polarization degree than any other accreting NSs in LMXBs with X-ray luminosities at least three orders of magnitude greater. In those systems, polarized X-ray emission is detected with polarization degrees of only a few percent and is primarily attributed to Comptonization, where seed photons from one or several regions near the NS scatter off high-energy particles ([119]; Appendix G). These regions are usually identified as areas on the NS surface where accreted material spreads due to rotational motion, transition zones where accreted material moves from the disk to the NS surface, or extended, geometrically thin plasma regions partially covering the disk. In some cases, additional polarized emission has been attributed to photons reflecting off the disk surface. Recently, X-ray radiation with a polarization degree of a few percent was detected from an accreting MSP during a bright outburst and attributed to moderately fan-beamed emission originating from small, heated spots on the NS surface [99]. The spectral properties of accreting NSs in LMXBs in spectral states similar to J1023 remain largely consistent across a wide range of X-ray luminosities, including levels comparable to those of J1023 [129, 83]. Hence, if similar emission mechanisms were at play in J1023, we would expect a polarization degree of a few percent at most. Our measured polarization degree exceeds this expectation, suggesting that these mechanisms are unlikely to be solely responsible for the observed polarization in J1023. This conclusion is further supported by theoretical considerations. Recent studies have demonstrated that emission from the spreading layer is expected to have a polarization degree not exceeding 1.5% [14] (see also [47]). Similarly, thermal Comptonization in accretion shocks above the surface of accreting MSPs results in a polarization degree of only a few percent at most [16].
4.1.2 Scattering in an Accretion Disk Wind?
Scattering of emission by outflows of highly ionized material launched from the accretion disk along the equatorial plane – commonly referred to as disk winds (e.g. [42]) – may provide an alternative source of polarization. Although there is currently no observational evidence for disk winds in J1023, recent work [93] suggests that even if such winds existed and significantly contributed to the X-ray emission from J1023, this mechanism alone cannot readily account for the high polarization degree at its known orbital inclination (; [115]). This conclusion holds true regardless of the wind opening angle, optical depth through the wind, or the nature and angular distribution of the illuminating X-ray emission, including scenarios with isotropic unpolarized sources, thermal radiation from the disk, or Comptonization [93].
4.1.3 A Rotation-powered Mechanism?
A few models have been developed to explain the X-ray to optical polarization properties in isolated pulsars, including the outer gap and two-pole caustic models, as well as current sheet models (for a review, see [60]). Assuming that the polarized emission from J1023 originates from mechanisms similar to those in isolated rotation-powered pulsars, we can then compare the polarization properties predicted by these models with those observed in J1023.
The outer gap (OG; [30, 31, 61, 117]) and two-pole caustic (TPC; [44]) models propose that the high-energy emission of pulsars originates from the outer magnetosphere, along the last open magnetic field lines inside the light cylinder radius. The emission mechanisms involve synchrotron radiation and curvature radiation from particles accelerated within vacuum gaps or extended regions of strong electric fields. Pulse profiles are characterized by caustics formed due to relativistic effects like aberration and light travel time delays. This results in double-peaked light curves, with peak separation depending on the magnetic inclination angle and the observer’s viewing angle. These models predict polarization properties such as rapid swings in the PA correlated with the pulses and dips in PD at the pulse peaks. The dips are attributed to depolarization from overlapping emissions originating from regions with varying magnetic field orientations. In the TPC model, rapid PA swings are expected through both peaks of the pulse profile. In contrast, the OG model predicts a significant PA swing mainly at the second peak, and the first peak may not show a strong PA swing unless the emission occurs near the light cylinder [43].
The PA in J1023 remains relatively constant across the pulse profile, at odds with the rapid PA swings predicted by the TPC model and, to some extent, the OG model. The PD remains relatively stable across the phase without significant dips at the pulse peaks, inconsistent with the expected PD dips due to depolarization from overlapping emissions.
Current sheet models [104, 28] focus on emission from particles accelerated in the equatorial current sheet beyond the light cylinder. Magnetic reconnection in the current sheet accelerates particles, leading to high-energy synchrotron emission. Pulses are formed when the observer’s line of sight crosses the current sheet, resulting in one or two pulses per rotation. These models predict PD values typically 15–30%, with PD dips at the pulse peaks due to depolarization effects. Significant PA swings are expected at the pulse peaks due to the reversal of the magnetic field polarity across the current sheet.
J1023 does not exhibit the sharp PA swings predicted by current sheet models or dips in the PD at the pulse peaks. The data suggest the emission region has a more uniform magnetic field orientation, inconsistent with the magnetic field reversals expected in current sheet models.
Overall, these discrepancies indicate that the emission mechanisms in J1023 differ from those assumed in standard models for isolated pulsars.
4.1.4 Pulsar Wind – Accretion Flow Interaction?
A scenario for the high mode of J1023 suggests the presence of an active rotation-powered pulsar, an accretion disk, and a compact jet. In this scenario, the observed X-ray, UV, and optical pulses originate from synchrotron emission at the surface of a shock region formed where the pulsar particle wind – modulated at the NS spin period – collides with the inner accretion flow [98, 122]. This interaction is thought to occur just beyond the light cylinder radius (the distance at which corotation with the pulsar becomes physically impossible), which is 80 km for J1023. The observed lag of 150 s between optical and X-ray pulses is suggested to originate from differences in the cooling timescales of synchrotron photons at these energies, given the expected magnetic field strength at the shock [98, 70]. A compact jet is launched, emitting partially self-absorbed synchrotron radiation from radio to infrared frequencies. Previous work shows that the shock emission contributes 5% of the optical flux and 90% of the soft X-ray flux, while the compact jet emission accounts for 2.5 in the optical band and 10% in soft X-rays [10].
Electrons spiraling along toroidal magnetic field lines around the jet axis emit synchrotron radiation, with the polarization oriented perpendicular to the field lines. Given this magnetic field geometry and the orbital inclination of J1023, calculations indicate that the expected polarization degree should not exceed 25% [123, 87, 77]. In the -band, this translates to a maximum polarization degree of 0.6%, which is lower than the average optical polarization observed for J1023. In the IXPE energy band, the limit is 3%, again below our measured value. Hence, based on these considerations, we rule out the compact jet as the main source of the observed polarization.
If the observed low-level optical polarization is due to synchrotron radiation from the shock region, and a single population of electrons accelerated at the shock front produces synchrotron radiation from optical to X-rays, we would expect the intrinsic polarization in the optical band to propagate to higher energies. In this scenario, we estimate an intrinsic polarization degree of 12–17% in the soft X-ray band [10]. Moreover, if the same region produces both optical and X-ray synchrotron radiation, their average polarization angles should align closely [10]. Our measurements are fully consistent with these predictions, providing striking evidence that the polarized emission in J1023 is driven by shock emission.
Figure 5 shows the SEDs of the high-mode flux, polarized flux, and pulsed flux of J1023 across optical and X-ray wavelengths. The SED of the polarized flux follows the same power law that describes the SED of the pulsed flux. Moreover, both the polarized and pulsed fluxes are consistent with each other. These results indicate that the polarized and pulsed emissions share a common origin, providing strong support for the scenario in which the synchrotron-emitting shock region is the primary source of both in J1023.
Figure 6 illustrates the geometry of the system during the high mode. The pulsar wind, structured in toroidal segments with alternating magnetic field polarities, collides with the accretion flow, imprinting its magnetic field pattern onto the shock surface. Particles accelerated in the region of the shock visible to the observer gyrate along these field lines, emitting synchrotron radiation predominantly in the direction perpendicular to the shock surface. In this configuration, the electric field vector of the emitted radiation oscillates along the plane of the shock surface. The observed average of the X-ray and optical photons – defined as the angle between the North-South direction on the sky plane (perpendicular to the line of sight towards the observer) and the orientation of the electric field vector projected onto the sky plane – is determined by the geometry of the shock region as it appears from the observer’s viewpoint. The measured consistent with 0∘ indicates that, on average, the electric field vectors of the emitted X-ray and optical photons align with the North-South axis on the plane of the sky. The toroidal symmetry of the magnetic field and the positioning of the emission region in the pulsar equatorial plane ensure that the projected direction of the electric field vector – and thus the polarization angle – remains constant from the observer’s viewpoint, regardless of the pulsar rotational phase.
In the scenario outlined above, the upper limit on the radio linear polarization fraction in the high mode is consistent with partially self-absorbed synchrotron radiation from the optically thick portion of the compact jet (for more details on this result and on the polarization measurements in the low mode, see Appendix F).
4.2 Synchrotron Emission across Scales
Our findings reveal striking similarities between J1023 and pulsar wind nebulae (PWNe) regarding emission mechanisms and magnetic field geometry. In PWNe, the high-energy emission is attributed to synchrotron radiation produced by electrons accelerated within highly ordered, predominantly toroidal magnetic fields. These fields are shaped by the pulsar wind as it wraps around the pulsar and is compressed within the nebula (see, e.g., [105]). However, the polarization degree observed in a few PWNe is higher than in J1023. For instance, the Vela PWN exhibits an average soft X-ray polarization degree of 45%, with inner regions reaching up to 70% [131]. Additionally, a polarization degree exceeding 60% has been detected in the outskirts (with 90% confidence; [84]). Similarly, the Crab PWN shows a spatially integrated PD of 20% [20], with some regions exhibiting polarization degrees between 45% and 50% [84, 20]. The PWN around the pulsar PSR B054069 reaches a PD of 25% [132].
The distinct geometries and scales of PWNe compared to J1023 may play a crucial role in the observed differences. In PWNe, the pulsar wind expands into the surrounding environment over large distances, with the wind termination shock occurring at 0.1 pc from the pulsar [64, 103]. This extended region facilitates the development of large-scale, toroidal magnetic fields with highly ordered structures. At the same time, it disperses any coherent pulsations emitted by the NS. Consequently, PWNe are not expected to produce pulsed emission through the same shock emission mechanism described for J1023. In contrast, J1023 features a much more compact emission region located 100 km from the pulsar. At such close proximity, the interaction between the pulsar wind and the inflowing material from the accretion disk is more disruptive to the striped wind structure due to the higher density and turbulence of the accretion flow. This interaction likely results in a more tangled magnetic field at the shock region, which may be the main factor accounting for the lower polarization degree observed in J1023 compared to PWNe.
5 Conclusions
We have conducted the first multiwavelength polarimetric campaign of the tMSP PSR J10230038, using observations across the X-ray, optical, and radio bands. The results of our analysis can be summarized as follows:
Timing analysis of the IXPE data revealed coherent X-ray pulsations at the NS spin period. The measured spin frequency suggests that the rotational evolution of the pulsar in its subluminous X-ray state has remained largely unchanged compared to its behavior during the radio pulsar state.
During the high mode, we detected polarized X-ray emission with a polarization degree of and angle of (1) in the 2–6 keV energy range. During the low mode, we derived an upper limit on the polarization degree of 26% at 90% c.l.. We do not observe any significant variability in the polarization properties along the NS rotational phases or the binary orbital phases.
Optical polarization measurements revealed a polarization degree of , with a polarization angle aligning closely with that of the X-ray emission. This strongly suggests a common origin for the polarized emission in both the optical and X-ray bands.
No polarized emission was detected at radio wavelengths, down to 3 upper limits on the polarization fraction of 9% in the high mode and 19% in the low mode.
The SEDs of the polarized and pulsed fluxes align with a single power-law model across optical and X-ray energies, indicating that the polarized emission originates from the same processes driving the pulsed emission.
The polarization degree in the high mode is higher than that seen in accreting NSs in LMXBs, indicating that standard accretion-powered emission mechanisms cannot fully explain the observed polarization. Additionally, scattering in an accretion disk wind alone is unlikely to account for the high polarization degree. Moreover, the lack of significant variability in the X-ray polarization properties across the pulsar rotational phases disfavors scenarios where the polarized signal arises as if the NS behaves like an isolated pulsar.
Our findings support a scenario where the polarized emission is produced by synchrotron radiation at a shock formed by the interaction between the pulsar wind and the inner accretion disk. This gives insights into how the pulsar wind interaction with its environment shapes the observed emission properties.
Our derived upper limits on the radio polarization fraction are consistent with the presence of a partially self-absorbed synchrotron-emitting compact jet during the high mode. During the low mode, the constraints align with expectations from synchrotron emission, considering potential depolarization effects.
These findings enhance our understanding of the complex interplay between accretion-powered and rotation-powered processes in tMSPs, and establish multiwavelength polarimetry as a crucial diagnostic tool for uncovering the physical mechanisms at work in these systems. Future observations, particularly those capable of measuring the polarization of pulsed emission, will be essential to further elucidate the nature of the emission processes in J1023 and similar objects.
Acknowledgments
We thank the NICER principal investigator, Keith Gendreau, for approving our Target of Opportunity (ToO) request and the operation team for executing the observations; we also thank the Swift deputy project scientist, Brad Cenko, and the Swift duty scientists and science planners, for making the Swift ToO observations possible. We thank Alexandra Veledina and Juri Poutanen for fruitful discussions on the contribution of synchrotron emission from jets to polarization in X-ray binaries. We thank Alessio Marino for useful discussions on the spectral properties of low-mass X-ray binaries.
This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). This research is based on observations collected at the European Southern Observatory under ESO programme 113.27RE. IXPE is a joint US and Italian mission. NICER is a 0.2–12 keV X-ray telescope operating on the International Space Station, funded by NASA. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.
M.C.B. acknowledges support from the INAF-Astrofit fellowship. F.C.Z. acknowledges support from a Ramón y Cajal fellowship (grant agreement RYC2021-030888-I). S.C. and P.D.’A. acknowledge support from ASI grant I/004/11/5. A.D.M. and F.L.M. are supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-INAF-2022-19-HH.0. A.P. and D.d.M. acknowledge financial support from the Italian Space Agency (ASI) and National Institute for Astrophysics (INAF) under agreements ASI-INAF I/037/12/0 and ASI-INAF n.2017- 14-H.0 and from INAF ‘Sostegno alla ricerca scientifica main streams dell‘INAF’, Presidential Decree 43/2018, from INAF ‘SKA/CTA projects’, Presidential Decree 70/2016. F.C.Z., S.Ca., P.D.’A., A.P., D.d.M. and G.I. acknowledge financial support from INAF-Fundamental research astrophysics project “Uncovering the optical beat of the fastest magnetised neutron stars” (FANS) and the Italian Ministry of University and Research (MUR) under PRIN 2020 grant No. 2020BRP57Z “Gravitational and Electromagnetic-wave Sources in the Universe with current and next-generation detectors (GEMS)”. AP acknowledges support from the Fondazione Cariplo/Cassa Depositi e Prestiti, grant no. 2023-2560. A.K.H. is supported by NSERC Discovery Grant RGPIN-2021-0400. D.M.R. and K.A. are supported by Tamkeen under the NYU Abu Dhabi Research Institute grant CASS. D.F.T. is supported by the grant PID2021-124581OB-I00 funded by MCIU/AEI/10.13039/501100011033 and 2021SGR00426. F.C. acknowledges support from the Royal Society through the Newton International Fellowship programme (NIF/R1/211296). G.I. is supported by the AASS Ph.D. joint research program between the University of Rome “Sapienza” and the University of Rome “Tor Vergata” with the collaboration of the National Institute of Astrophysics (INAF). N.R. is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Consolidator Grant “MAGNESIA” No. 817661) and the Proof of Concept “DeepSpacePulse” (No. 101189496). F.C.Z. and N.R. acknowledge support from grant SGR2021-01269 from the Catalan Government. This work was also supported by the Spanish program Unidad de Excelencia María de Maeztu CEX2020-001058-M and by MCIU with funding from European Union NextGeneration EU (PRTR-C17.I1).
Softwares
ASTROPY v.7.0, a community-developed core Python package for Astronomy [7]; CASA v.6.5 [27]; DAOPHOT [114]; DUST-EXTINCTION v.1.6 [56]; HEASOFT v.6.34 [92]; IPYTHON v.8.30 [102]; IRAF v.2.18, distributed by the National Optical Astronomy Observatory, which is operated by the Association of Universities for Research in Astronomy, Inc., under a cooperative agreement with the National Science Foundation (https://iraf-community.github.io/); IXPEOBSSIM v.31.0.3 [11], documented at https://ixpeobssim.readthedocs.io/en/latest/; MATPLOTLIB v.3.9.3, a Python library for publication-quality graphics [69]; NICERDAS v.13; NUMPY v.1.24.0 [62]; PANDAS v.2.2.2 [128]; PINT v.1.1 [86]; QUARTICAL v.0.2.3 [75]; SAOImageDS9 v.8.6, a tool for data visualization supported by the Chandra X-ray Science Center (CXC) and the High Energy Astrophysics Science Archive Center (HEASARC) with support from the James Webb Space Telescope Mission office at the Space Telescope Science Institute for 3D visualization [73]; SCIPY v.1.14.1 [125]; Swift Reduction Package (SRPAstro) v.4.9.0 (https://pypi.org/project/SRPAstro/); WSClean v3.4 [66]; XRONOS v.5.22 [113]; XSPEC v.12.14.0i [6].
Appendix
Telescope/Instr. | Obs./Prg. Id | Setup | Start – End time | Exposure | Band |
YYYY Mmm DD hh:mm:ss (UTC) | (ks) | ||||
IXPE/DUs | 03005599 | 2024 May 29 12:30:51 – 2024 Jun 14 14:57:45 | 675 | 2–8 keV | |
NICER/XTI | 7034060107 | 2024 Jun 01 02:42:20 – 2024 Jun 01 21:36:39 | 4.3 | 0.5–10 keV | |
Swift/XRT/UVOT | 00033012239 | PC / Event mode | 2024 Jun 04 22:14:59 – 2024 Jun 04 23:59:54 | 1.5 | 0.3–10 keV / |
Swift/XRT/UVOT | 00033012240 | PC / Event mode | 2024 Jun 05 01:23:09 – 2024 Jun 05 01:48:53 | 1.5 | 0.3–10 keV / |
VLA | 24A-476 | B configuration | 2024 Jun 04 23:10:31 – 2024 Jun 05 00:36:20 | 5.1 | |
VLT/FORS2 | 113.27RE.001 | Woll. prism + HWP | 2024 Jun 04 23:34:37 – 2024 Jun 05 00:39:23 | 3.9 | |
Swift/XRT/UVOT | 00033012241 | PC / Event mode | 2024 Jun 06 02:39:12 – 2024 Jun 06 04:48:54 | 2.2 | 0.3–10 keV / |
NICER/XTI | 7034060108 | 2024 Jun 08 18:53:00 – 2024 Jun 08 22:17:30 | 1.4 | 0.5–10 keV |
Appendix A Observations & data processing
The multiband data presented in this work were acquired by different space-borne and ground-based telescopes (Table 2). In the following, we describe the observations and the procedures adopted to process these data. In the following analyses, all timestamps are referenced to the barycenter of the Solar System using the Jet Propulsion Laboratory planetary development ephemeris DE440 [100] and the position derived from observations with the Very Long Baseline Array [36]: RA = 10h23m47s69 and decl. = +00∘38′40.′′8 (Julian epoch date J2000.0; the uncertainties in both coordinates are 2 as).
A.1 IXPE
The Imaging X-ray Polarimetry Explorer (IXPE), a joint mission between the National Aeronautics and Space Administration (NASA) and the Italian Space Agency (ASI), was launched on December 9, 2021. The observatory consists of three identical telescopes, each equipped with an X-ray mirror module assembly that uses grazing-incidence optics to focus X-rays onto a detector unit (DU). Each DU contains a polarization-sensitive gas-pixel detector (GPD) that measures linear X-ray polarization by imaging the tracks of photoelectrons produced by incoming X-rays absorbed in the special fill gas [12, 39]. The GPDs are capable of recording the arrival time of X-ray photons with a time resolution of 100 s (for information on IXPE, see [111, 127]).
IXPE observed J1023 during two sessions: the first from 2024 May 29 at 12:32 UTC to 2024 June 2 at 18:12 UTC (MJD 60459.52 – 60463.76), and the second from 2024 June 3 at 21:15 UTC to 2024 June 14 at 14:58 UTC (MJD 60464.88 – 60475.62). The total exposure time per DU was 675 ks. Data processing was performed using the ixpeobssim package [11]. Both spectral and spectro-polarimetric analyses were conducted with HEASoft and the standard FTOOLS, utilizing the latest files in the IXPE Calibration Database (XRT version released on 2024 August 20, GPD version released on 2024 February 28).
Instrumental background in the photon event lists was reduced by applying the rejection algorithm by [40]777https://github.com/aledimarco/IXPE-background. Given the source faintness, background subtraction was also performed. Source and background photons were selected from the cleaned level-2 event data. Source photons were selected from a circular region with a 60′′ radius centered on the source, corresponding to roughly the 95% encircled energy fraction for the point spread function of an on-axis source. The background was selected from an annular region with inner and outer radii of 150′′ and 300′′, respectively (Fig. 7). Even after background rejection, the background signal dominated above 6 keV. Consequently, the polarimetric analysis was restricted to the 2–6 keV energy band.
A.2 NICER
The Neutron Star Interior Composition Explorer Mission (NICER; [52]) is a non-imaging X-ray observatory on board the International Space Station, with a field of view of 30 arcmin. The X-ray timing instrument of NICER consists of 56 co-aligned X-ray concentrator optics paired with silicon-drift detectors, 52 of which are currently operational. NICER is sensitive to photon energies in the 0.2–12 keV range, provides an effective collecting area of 1900 cm2 at 1.5 keV, and offers a time resolution below 100 ns. During the IXPE campaign, we conducted a couple of NICER observations of J1023 as part of the director’s discretionary time (see Table 2 for a journal of the observations). These observations were aimed at guiding the selection of high and low-mode episodes in the IXPE dataset and were performed during orbit night-time, ensuring high-quality data unaffected by the optical light leak caused by a damage of the thermal shield on one detector sustained in May 2023888See https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/light-leak-overview/.. We reprocessed the data using the NICER Data Analysis Software (NICERDAS) version 13 distributed along with the HEASoft software package and the calibration products released on 2024 February 6. We calibrated and screened the raw unfiltered data and created level-2 cleaned event files using the nicerl2 pipeline with default settings. We extracted time series over the 0.5–10 keV energy range, binned at 10 s, using the nicerl3-lc pipeline. To estimate the background level, we applied the SCORPEON model999https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/scorpeon-overview/ along with the latest geomagnetic data available at the time of analysis (see the yellow dashed lines in Fig. 9, bottom panel).
A.3 Swift
We performed a couple of observations with the Neil Gehrels Swift Observatory (Swift; [51]) simultaneously with the VLA and VLT observations on 2024 June 4–5 (Table 2). During these observations, Swift utilized the X-ray telescope (XRT; [21]) in photon counting mode (readout time of 2.51 s) and the Ultra-Violet/Optical Telescope (UVOT; [106]) in event mode (readout time of 11 ms) with the UVM2 filter in place (central wavelength of 226 nm; full width at half maximum of 52.7 nm). Data were processed using the calibration files released on 2024 May 22 for XRT and on 2024 February 1 for UVOT.
After performing a standard screening of the XRT data, we collected source photons using a circular aperture with a radius of 47.2′′ centered on J1023 and estimated the background level using an annular region with inner and outer radii of 94.4′′ and 188.8′′, respectively. We then extracted background-subtracted time series in the 0.5–10 keV range, binned at 50 s (Fig. 9).
For the UVOT data, we used the coordinator and uvotscreen tools to convert raw coordinates to detector and sky coordinates, filter out hot pixels on the charge-coupled devices, and obtain cleaned event lists. We adopted a circular extraction region of radius 5′′ centered on the position of J1023 and a circular extraction region of radius 10′′ away from the source to collect source and background photons, respectively. We ensured that the extraction region was always centered on the source position in each observation snapshot, and extracted background-subtracted time series binned at 30 s using uvotevtlc (Fig. 9).
A.4 VLT
We observed J1023 with the FOcal Reducer/low dispersion Spectrograph 2 (FORS2; [2]) mounted on the Unit Telescope 1 (Antu) of the 8.2-m Very Large Telescope (VLT) at Paranal Observatory (Chile) in polarimetric mode (Table 2). The observations began on 2024 June 4 at 23:34:37 UTC (MJD 60465.982) and ended on 2024 June 5 at 00:39:23 UTC (MJD 60466.027), and were performed under photometric conditions (seeing 0.3”). A total of 56 images were acquired using the optical filter (; central wavelength of 655 nm; full width at half-maximum of 165 nm), each with an exposure of 20 s, resulting in a total observing time of 3946 s, overheads included. A Wollaston prism was placed in the light path of the instrument, splitting the incident radiation into two orthogonally polarized beams (ordinary and extraordinary). A Wollaston mask was used to prevent these beams from overlapping on the CCD. Moreover, a rotating half-wave plate (HWP) was installed, allowing images to be captured at four distinct angles () relative to the telescope axis: . This step is critical to obtain polarization measurements because combining images taken at multiple (more than 2) angles is necessary to determine the level of linear polarization. Four sets of images were acquired with this setup, one for each angle of the HWP. This process was repeated 14 times to obtain the entire dataset.
Images were processed by subtracting an average bias frame and dividing by a normalized flat field. Aperture photometry was performed using the daophot tool [114], using a 6-pixel aperture. Using Equations 1–2 of [9], we calculated the normalized Stokes parameters and for the linear polarization. These parameters have not been corrected for instrumental contributions to the linear polarization. However, FORS2 routinely observes unpolarized standard stars to monitor instrumental polarization, and this has consistently remained very low (below 0.3%) across all bands over the past decade.
To estimate the degree and angle of linear polarization, we evaluated the parameter for each of the HWP angles, following [41] and references therein (see also [34, 9, 10]). describes the linear polarization along the direction (e.g., ), and it is defined as:
Here and are the extraordinary and ordinary fluxes at a given angle , respectively, and and are the corresponding fluxes for a non-polarized standard star in the field. The relationship between , the polarization degree and the angle is given by:
The parameters and were determined through a two-step process: 1) maximizing the Gaussian likelihood function to provide an initial estimate for and ; 2) refining the estimates from 1) using a Markov Chain Monte Carlo (MCMC) procedure [67] (for further details on the algorithm, see [9, 10]). The best-fit values were obtained as the median of the marginal posterior distributions from the MCMC procedure. The uncertainties were estimated using the 16th to 84th percentiles of these distributions.
Since is calculated by normalizing with a non-polarized star in the field, the resulting value for the polarization degree obtained after fitting the curve is already corrected for instrumental effects. However, the polarization angle still requires correction using the Stokes parameters of a polarized standard star. To achieve this correction, we observed the star Vela 1-95 and measured its Stokes parameters to accurately determine its polarization angle. This measured angle was then compared to the tabulated value, resulting in a correction of applied to the polarization angle of J1023. Finally, to correct for the interstellar contribution, instead of using only one non-polarized star, we used a group of reference stars in the same field as the target in Equation 2, assuming these stars are non-polarized. This method allows the polarization degree of the target to be corrected for both instrumental and interstellar contributions. It is crucial to ensure that the chosen reference stars all exhibit the same polarization degree, which would be a good indicator of their intrinsic non-polarized nature (see e.g., [9]). Since the field of J1023 is relatively empty of optical sources, we could select only four comparison stars. However, these stars clustered very well together in the plane, allowing us to determine the instrumental and interstellar polarization quite precisely in each image. We note that, as reported by [10], the maximum interstellar contribution to the linear polarization degree of J1023 is quite low (), and comparable to the polarization measured for the instrument configuration.
A.5 VLA
We acquired radio observations of J1023 – simultaneous with the IXPE and VLT monitoring – using the Karl G. Jansky Very Large Array (VLA; Project ID 24A-476) radio interferometer. Our full observations, including both target and calibrator scans, began on 2024 June 4 at 23:00:00 UTC (MJD 60465.958) and ended on 2024 June 5 at 02:59:20 UTC (MJD 60466.125). The source underwent a bright multi-wavelength flare starting at 00:36:00 UTC on 2024 June 5 (MJD 60466.025); in this work, we only include the analysis of the pre-flare dataset (i.e., 2024 June 4 at 23:00:00 to 2024 June 5 at 00:36:00, with 1.2 hr on source).
We observed in C-band (central frequency of 6 GHz) using the 3-bit sampler, which breaks the total un-flagged bandwidth of 4096 MHz into 32 spectral windows, each comprised of 64 2 MHz frequency channels. We used 3C286 as the flux and bandpass calibrator, J10240052 as the nearby complex gain calibrator (i.e., the secondary), and the unpolarized source, J14072827, as the leakage calibrator. The target monitoring of J1023 was broken into 9-min scans, with two 1-min secondary scans bookended each target scan.
We flagged and calibrated the parallel-hand visibilities with the Common Astronomy Software Applications package (CASA) VLA pipeline [26]. After the pipeline processing was completed, we inspected the visibilities, removing residual radio frequency interference (RFI) through a combination of manual editing and CASA auto-flagging algorithms (i.e., rflag and tfcrop). After the parallel-hand calibration (which is only sufficient for Stokes ) we calibrated the cross-hand visibilities following the procedure outlined in the VLA tutorials101010The tutorial can be found here: https://casaguides.nrao.edu/index.php/CASA_Guides:Polarization_Calibration_based_on_CASA_pipeline_standard_reduction:_The_radio_galaxy_3C75-CASA6.5.4.
We created Stokes images with the WSClean imager [94]; WSClean is a CLEAN-based deconvolution algorithm [66] that removes sampling artefacts from interferometric images. To produce our final data products, we first performed a shallow unmasked deconvolution of the full 10′ field of view (FOV), i.e., we created preliminary images. We made a deconvolution mask using the unmasked Stokes image and, with the mask, re-imaged the target field. Adopting the masked image as a model, we performed phase-only self-calibration on the target visibilities with the software package QuartiCal [75]. We then refined the mask and repeated the imaging and self-calibration procedure a second time. In the FOV of J1023, there exists, among other fainter background sources, a bright radio Galaxy (J102358.2003826) at an angular offset of . Using the self-calibrated model, we subtract out the visibilities for the background sources, allowing for narrow FOV images that are minimally affected by aliasing.
During the VLA observations, J1023 had an average flux density of 240 Jy at 6 GHz. We created two sets of images using our background-subtracted visibilities. First, we broke the observations into 1-minute imaging intervals, allowing us to (temporally) resolve the radio variability. The 1-minute images revealed that J1023 underwent a mini-flare lasting 8 min and reaching a maximum flux density of 500 Jy. We created four additional images that combined multiple target scans to get deeper polarization limits, sacrificing temporal resolution for higher signal-to-noise. The deep imaging time intervals were 23:10:31–23:52:25 UTC (June 4), 23:52:25–00:00:22 UTC (June 4 to June 5) and 00:00:22.5–00:36:20 UTC (June 5); these intervals correspond to the times before the mini-flare, the mini-flare, and the times after the mini-flare. The fourth image combined the times before and after the mini-flare.
In each imaging interval, WSClean outputted 16 channelized images (evenly spaced in frequency) and a single frequency-averaged ‘Multi-Frequency-Synthesis’ (MFS) image. We measure the target’s Stokes flux density from the MFS image, using the CASA task imfit. imfit models the intensity distributions as an elliptical Gaussian; given that J1023 is a point source, we enforce a synthesized beam-shaped Gaussian fit. We quantified the uncertainty on the flux density with the root-mean-square (RMS) noise extracted from a nearby source-free region. Using the CASA task imstat, we measure the RMS using a circular aperture with an area equal to 100 synthesized beam area. Additionally, for MFS images where J1023 was detected at (corresponding to in each channelized image), we calculate the intra-band power-law spectral index from channelized flux densities (using scipy.curvefit) , which describes how the flux density varies with observing frequency as . We found that remains nearly constant during the high-mode episodes that precede (H1) and follow (H2) the low mode, but decreases during the low mode (L). Specifically, we obtain , and .
Lastly, we produced linear polarization intensity images for each imaging interval but did not detect any polarized emission. As a result, we calculate 3 ( 99.7%) percent upper-limits from the -images following the prescription by [120]. We obtain , and . By stacking high-mode data, we obtain .
Appendix B Selection of Emission Modes in the IXPE Dataset
The top panel of Fig. 9 shows the time series for the IXPE observations, whereas the middle and bottom panels show the six time intervals where there was overlap between IXPE and NICER observations, totaling 2 ks. In each interval, the average background intensity in the NICER dataset was 10–15% of the total emission (source plus background), except for the last time segment, where the emission was consistently faint, and the background intensity accounted for 30% of the total emission. Nonetheless, during these overlapping time intervals, we did not observe any significant variation in the background level possibly correlated with mode-switching episodes, except for the fifth interval, where a slightly enhanced background level was detected at the exit of a low mode. This increase in the count rate should be approached with caution. Overall, we detected four low modes in the NICER data that were covered by IXPE: one in the second time segment, one in the third, one in the fifth, and one throughout the entire duration of the sixth segment.
To select the appropriate thresholds for the count rates in order to select the distinct emission modes in the IXPE data, we proceeded as follows. Over the past decade, the net count rate observed in all previous observations of J1023 carried out by the European Photon Imaging Cameras (EPIC) on board XMM-Newton decreased by a factor of approximately six during the low mode compared to the average net count rate (e.g., [18, 33]). The average count rate of J1023 measured by IXPE after combining data from the three DUs is 0.09 counts s-1, with about 30% of this rate contributed by the background level. Consequently, during low-mode episodes, typically only 0 or 1 net photon from the source, plus 3–4 photons from the background, would be detected by IXPE in 100-s time bins. In fact, during the low modes covered by NICER, the highest count rate measured in the IXPE time series binned at 100 s is 0.05 counts s-1 (summing up the source and background contributions). We then classified time intervals in the IXPE light curves as low modes whenever the count rate in the 100-s binned time series was less than or equal to 0.05 counts s-1. This criterion resulted in the selection of about 182 ks of low-mode data, which corresponds to about 27% of the total exposure time. This value is consistent with the fraction of time that J1023 was detected in the low mode in previous observations [5, 18, 33].
To differentiate between high and flaring modes, we extracted the distribution of count rates above 0.05 counts s-1 and modeled it using two Gaussian functions. We then classified a 100-s time bin as a flare mode if its count rate was at least 4 standard deviations above the mean count rate of the Gaussian distribution representing the high mode, which corresponds to a threshold of 0.2 counts s-1 (Fig. 10). This approach resulted in the selection of 20 ks of flaring-mode data, which represents 3% of the total duration of the IXPE observations. This percentage is again consistent with the fraction observed in previous observations [5, 18, 33].
Appendix C Timing Analysis of the IXPE Dataset
Previous studies have shown that coherent X-ray pulsations at the NS spin period in J1023 are detected only during the high mode [5, 71, 98, 70]. Hence, to search for a pulsed signal in the IXPE data, we retained only source photons collected during the high mode, and restricted the analysis to photons in the energy range 2–6 keV. We corrected the photons’ arrival times for the motion of the pulsar along the binary orbit, setting the orbital period and the projected semi-major axis of the circular orbit to the values reported by [3], = 17115.5216592 s and asini/c = 0.343356 light-seconds (the uncertainties in these parameters are of the order 10-7 s and 10-6 light-seconds, respectively). To measure the epoch of passage of the pulsar at the ascending node of the orbit (), we performed an epoch folding search on the IXPE dataset around the spin frequency value expected according to the previously measured secular evolution [71, 22], sampling each spin cycle in 16 phase bins. We used a grid of values centered on the value predicted according to the long-term evolution measured by [70], MJD, with a spacing of 0.2 s. We recovered the coherent signal with an epoch folding variance of (corresponding to a single-trial false alarm rate of ) at a spin frequency of Hz and = 60461.979580(14) MJD. The uncertainty on the spin frequency was evaluated following [80], while that on was taken as the half-width at half-maximum of the Gaussian fit to the pulse variance distribution. The timing model, compatible with the Tempo2 [65] and PINT [86] pulsar timing packages, will be made available at Zenodo.
The top panel of Fig. 13 shows the IXPE background-subtracted double-peaked pulse profile. We calculate the fractional RMS variability amplitude () of the pulsed signal following the method by [121] and obtain = 8.1%0.8%. This value is consistent within the uncertainties with those derived in previous studies [5, 71, 98].
We obtained compatible results by analysing 0.5–6 keV NICER data in the high mode taken on 2024 June 1. The timing parameters had slightly larger uncertainties due to the lower number of photons recorded () compared to IXPE () and the uneven coverage.
The spin frequency measured from IXPE data is compatible within the uncertainties with the value predicted from the quasi-coherent timing solution derived from optical data by [22] ( Hz), who found a spin down rate very close to the value observed when J1023 was detected as a radio pulsar before the state transition occurred in 2013. On the other hand, it differs by 16090 nHz from the value predicted by [71] ( Hz), who reported a spin-down rate about 20% faster (Fig. 11). Even though this difference is still compatible within a 3 c.l., our measurement tends to favor a scenario in which the rotational evolution of the pulsar in the subluminous X-ray state has remained largely unchanged compared to its behavior during the radio pulsar state.
Dataset and model summary | |
Pulsar name | J1023+0038 |
MJD range | 60459.5257—60475.6217 |
Data span (d) | 16.096 |
Solar system ephemeris | DE440 |
Time unit | TDB |
Parameters | |
POSEPOCH, Reference epoch for position () | 55000.00 |
RAJ, Right Ascension (J2000) | 10h23m47.687198s |
DECJ, Declination (J2000) | +00∘38′40.84551′′ |
F0, Spin frequency () | 592.42146705(9) |
PB, Orbital period () | 0.1980963155 |
A1, Projected semi-major axis of pulsar orbit (-) | 0.343356 |
TASC, Epoch of ascending node () | 60461.979580(14) |
Appendix D X-ray Polarimetric Properties
D.1 Model-independent Analysis
We first perform a model-independent polarimetric analysis of the IXPE data using the formalism outlined by [76] implemented in the ixpeobssim package [11] under the pcube algorithm in the xpbin routine. We compute the normalized, background-subtracted Stokes parameters and , where , and are the Stokes parameters. From these, we derive the polarization degree and polarization angle (also referred to as the electric vector position angle) , measured counterclockwise from the North celestial pole toward the East. Hereafter, uncertainties are given at the 68.27% c.l. unless noted otherwise.
In the 2–6 keV energy band, we obtain an MDP of 11% at a 99% c.l. and %. The polarization in the 2–3 keV energy range is consistent with zero within , while low-significance polarization is consistently detected across the 3–4, 4–5, and 5–6 keV energy ranges (Fig. 2). In the 3–6 keV range, we find an MDP of 15% and measure a polarization degree that exceeds this threshold, %, with . To assess the significance of this detection, we use the ratio of the polarization degree to its standard deviation, , as the primary indicator of statistical significance111111https://heasarc.gsfc.nasa.gov/docs/ixpe/analysis/IXPE_Stats-Advice.pdf. Using the distribution for two dofs, we calculate a confidence level (c.l.) of 99.4%.
Applying the same analysis to high-mode data, we obtain an MDP of 12% in the 2–6 keV range, with % and . The energy-resolved analysis corroborates the results from the entire data set (Fig. 3). In the 3–6 keV range, we find an MDP of 16% and %, with .
For low-mode and flaring-mode data, the significance was too low for a reliable polarimetric analysis. For the low mode, we obtained an MDP of 50% and % at a 90% c.l.. For the flaring mode, we derived an MDP of 30% and % at a 90% c.l..
D.2 Spectro-polarimetric Analysis
To enhance the significance of the polarimetric measurements, we conducted a spectro-polarimetric analysis of the data as outlined by [116], following the weighted analysis method prescribed by [38], where each photoelectron track recorded by the DUs is assigned a weight based on its ellipticity. We extracted the , and spectra for each DU using the xpbin tool’s PHA1 algorithms in ixpeobbsim, applying a constant energy binning of 200 eV. We assigned the response matrices 20240701-alpha075 available in the calibration database and computed modulation response functions and ancillary response files using ixpecalcarf.
Previous observations have shown that the X-ray spectrum of J1023 during its high mode is well described by an absorbed power-law model (e.g., [18, 33, 24]). Hence, the adopted model includes a power-law component corrected for interstellar medium effects using the Tübingen-Boulder model [130] with photoelectric cross-sections by [124], convolved with an energy-independent polarization model. To account for potential discrepancies between the effective areas of the different DUs, we also included a cross-normalization constant, which was fixed at unity for DU1 to serve as the reference, and allowed to vary for DU2 and DU3. Thus, the final model employed within the XSpec spectral fitting package [6] is const TBabs powerlaw polconst. We fit this model simultaneously to the Stokes energy spectra of all DUs in the 2–6, 2–3 and 3–6 keV energy ranges using the forward-folding, maximum likelihood procedure implemented in XSpec.
Observations over the past decade have shown that the X-ray spectral shape of J1023 has remained remarkably stable throughout this time span [18, 33, 24]. Hence, during the fitting process, we fixed the absorption column density () and power-law photon index () at the values measured in previous works: cm-2, . All other parameters were allowed to vary. The results are summarized in Table 1 and are consistent with those derived from the model-independent analysis. Figure 4 shows the confidence contours for the measurements of and obtained over distinct energy bands, drawn using the steppar command in XSpec.
It is noteworthy that previous analyses of the distributions of mode durations extracted from observations carried out by XMM-Newton EPIC [5] show that 40% of low-mode episodes last less than 100 s. This implies that our method for selecting the emission modes inherently misses short low-mode episodes totaling 10% of the total duration of the IXPE observations. According to [10], the emission during the low mode is expected to be polarized to a lower extent compared to the high mode due to the much smaller contribution of the shock emission. Even in the most extreme case where the low-mode emission is fully unpolarized, the X-ray polarization degree during the high mode selected using our criteria would be enhanced by a factor of [40]. Here, and represent the background-subtracted count rates in the IXPE band in the low and high modes, respectively. This small increase, equivalent to about 0.2% in the polarization degree, is within the margin of the measurements uncertainties, meaning that as a matter of fact the intrinsic X-ray polarization degree of the high-mode emission remains unchanged within the limits of our observations.
D.3 Phase-Resolved Analyses
We conducted both rotational and orbital phase-resolved analyses of the X-ray polarimetric properties. Using the photonphase tool in the PINT software package [86], we computed the rotational and orbital phases of each photon collected in the high mode in the 2–6 keV energy range by IXPE, based on the timing model derived from the IXPE data analysis. For each analysis, we split the data into 10 phase bins, and calculated the Stokes parameters for each bin using the pcube algorithm within ixpeobssim, following the same model-independent procedure described above and applying background subtraction.
Figure 12 shows the Stokes parameters in distinct pulsar rotational phases (left) and binary orbital phases (right). The left panel of Figure 13 displays the Stokes parameters, and PAX,H as a function of the rotational phase. We do not measure statistically significant (anti-)correlations between the normalized rate and the Stokes parameters. Specifically, Spearman rank and Kendall tau correlation analyses between the normalized rate and yielded coefficients of -0.25 (p-value = 0.49) and -0.16 (p-value = 0.60), respectively. Similarly, correlation analyses between the normalized rate and resulted in coefficients of -0.47 (p-value = 0.17) and -0.33 (p-value = 0.22), respectively. Constant fits to the and values produced values of 8.5 and 11.8 for 9 dof, respectively. The right panel of Figure 13 shows the Stokes parameters, and PAX,H as a function of the orbital phase. Similarly, no significant variability was observed in and , with constant fits yielding values of 5.6 and 9.8 for 9 dof, respectively.
Appendix E Optical Properties
All VLT observations resulted in significant detections (above 3), with a mean linear polarization level of (1). This value aligns with previous results obtained in the same band [8, 10], although it is slightly higher. The polarization angle is well-constrained, with a mean value of (1; measured after correcting for the polarization angle of the standard star). This value is consistent with that reported by [10], although our measurements are significantly more precise.
In addition to the polarimetric analysis, it is also possible to extract and calibrate fluxes to build a light curve. If we assume the flux loss is negligible, the sum of the intensities measured in both the ordinary and extraordinary beams for each image gives the total flux from the target. Therefore, we summed the ordinary and extraordinary fluxes for each image of J1023, as well as for two isolated field stars with R2 magnitudes listed in the USNO B1 catalogue [91]. The resulting flux was then calibrated against the flux of the reference stars.
Figure 1 shows the time evolution of the magnitude, Popt and PAopt. To quantify the intrinsic variability amplitude of the light curve, we computed the excess variance , where is the standard deviation of the magnitudes and is the mean square measurement error in magnitudes. We obtain ExcessVar 0.02 mag2, indicating moderate variability.
During the low mode episode, we observe an increase in the optical emission by a factor of 1.3, which contrasts with the behavior seen in the X-ray and UV bands. The origin of this optical increase is unclear, but it may be related to the broadening or enhanced prominence of hydrogen H and/or three helium He I emission lines. These lines, possibly originating from regions within the accretion disk or from mass outflows, fall within the wavelength range of the filter of FORS2 (572.5–737.5 nm). A recent high-time-resolution optical spectroscopic study of J1023 using the OSIRIS spectrograph mounted on the GranTeCan reported significant variability in the equivalent width and full width at half maximum of these lines, such as changes by a factor of two over minute timescales for the H line, with no clear correlation with the flux of the optical continuum [89] (see also [110, 59] for further evidence of highly variable emission line profiles). In contrast, no prominent emission lines are expected within the wavelength range of the UVM2 filter of the Swift UVOT, as shown in the UV spectrum by [63]. We emphasize that if the optical emission lines originate from an outflow, this observation would lend support to scenarios suggesting stronger outflows during the low mode [122, 10], which also result in increased radio emission. However, since similar variability in the optical light curve is also observed in other instances during the X-ray high mode, it is not possible to draw a definitive conclusion.
Epoch | Qopt | Uopt | Popt | PAopt |
(BMJD) | (%) | (∘) | ||
60465.9887 | 0.0131 0.0014 | -0.0009 0.0014 | 1.43 0.13 | 2.0 2.8 |
60465.9915 | 0.0127 0.0014 | 0.0009 0.0014 | 1.44 0.13 | -0.1 2.7 |
60465.9943 | 0.0101 0.0014 | 0.0005 0.0014 | 1.01 0.13 | 2.8 3.8 |
60465.9971 | 0.0074 0.0014 | -0.0037 0.0014 | 0.97 0.13 | -7.6 4.0 |
60465.9999 | 0.0129 0.0014 | -0.0012 0.0014 | 1.51 0.14 | -2.4 2.6 |
60466.0027 | 0.0090 0.0014 | -0.0044 0.0012 | 1.03 0.13 | -4.7 3.2 |
60466.0055 | 0.0131 0.0014 | -0.0069 0.0013 | 2.12 0.13 | -12.3 1.9 |
60466.0083 | 0.0113 0.0014 | -0.0041 0.0014 | 1.60 0.13 | -11.4 2.1 |
60466.0112 | 0.0101 0.0014 | -0.0018 0.0014 | 1.12 0.14 | -0.3 3.5 |
60466.0140 | 0.0129 0.0014 | 0.0021 0.0014 | 1.57 0.13 | 5.9 2.5 |
60466.0168 | 0.0138 0.0014 | -0.0028 0.0014 | 1.64 0.14 | -4.9 2.4 |
60466.0196 | 0.0152 0.0014 | -0.0030 0.0014 | 1.57 0.14 | 0.3 2.5 |
60466.0225 | 0.0108 0.0014 | 0.0023 0.0014 | 1.11 0.14 | -2.5 3.5 |
60466.0253 | 0.0106 0.0014 | -0.0037 0.0014 | 1.19 0.13 | -7.3 3.3 |
Popt exhibits moderate time variability, with a Median Absolute Deviation (MAD) of 0.23% and a normalized Excess Variance (NEV) of 0.05, indicating some intrinsic fluctuations. Its skewness is 0.58, suggesting a slight right-skewed distribution and indicating that deviations above the mean are more frequent. On the other hand, PAopt shows greater variability, with a MAD of 3.6∘, a significantly higher NEV of 12.7, and a skewness of -0.23, pointing to a slight left-skewed distribution where deviations below the mean are more common. The opposite trends in the skewness of Popt and PAopt are primarily due to their anticorrelated variability during the second peak of optical intensity (Figures 1 and 14). We measure and for the high mode; and for the low mode. Consistent with previous findings by [10], these measurements suggest a slight decrease in polarization degree when switching from the high to the low mode, while the polarization angle shows no significant variation within the uncertainties.
Figure 14 shows the time series of the Qopt and Uopt Stokes parameters. The two parameters exhibit distinct variability patterns. Qopt exhibits relatively low variability, with a MAD of 0.0013 and a NEV of 0.03, indicating minimal intrinsic fluctuation. Its skewness is -0.33, suggesting a slightly left-skewed distribution with more frequent deviations below the mean. Uopt shows more pronounced variability, with a MAD of 0.0016 and a significantly higher NEV of 1.8, indicating substantial intrinsic fluctuations. The skewness is 0.04, indicating an almost symmetric distribution, with a very slight tendency for deviations above the mean. These results suggest that Uopt is more variable and less symmetric than Qopt.
Figure 14 also shows the and Stokes parameters of J1023 on the plane for the whole dataset, including high modes (filled dots) and the low mode (filled diamond). The data points are color-coded according to the time they were recorded. Despite some intrinsic variation, the and values cluster around a common point, and there is no clear distinction between the high and low modes.
Appendix F Polarization Properties in the Low Mode; Constraints on Polarized Radio Emission in the two Modes
In the scenario outlined in Section 4.1.4, the upper limit on the radio linear polarization fraction in the high mode is consistent with partially self-absorbed synchrotron radiation from the optically thick portion of the compact jet. The exact cause of the switch to the low mode remains uncertain, though the detection of short millimeter-band flares during a few of these switches in past observations points to sudden ejections of magnetized plasma, likely originating from the innermost regions of the accretion flow. The ejection of this matter could thus force the shock region to move outward, reducing the X-ray flux and the pulsed emission. As the plasma propagates downstream in the compact jet, it becomes optically thin in the radio band [10]. This framework also accounts for the optical polarization properties measured during the low mode. If the shock moves outward, it may still contribute a measurable amount of polarized light. Moreover, the optical polarization angle observed in the low mode is consistent with that of the high mode within uncertainties, suggesting a common underlying emission mechanism for the optical linear polarization across both modes and reinforcing the idea that residual shock emission remains active in the low mode. In the soft X-ray band, where the jet could dominate up to 100% of the emission [10], the maximum expected polarization would be 25%. The upper limit for the X-ray polarization degree during the low mode, % at a 90% c.l., makes it impossible to draw firm conclusions from this measurement. Alternatively, the switch to the low mode may result from matter occasionally penetrating the light cylinder, temporarily forcing the system into a regime where the rapidly rotating magnetic field anchored to the NS propels the inflowing matter away. As matter expands outward, it becomes increasingly rarefied and more transparent to synchrotron emission at radio wavelengths [122].
At radio wavelengths, synchrotron radiation is expected to produce linear polarization fractions of up to 10% for optically thick emission and up to 70% for optically thin emission, assuming uniform magnetic fields and a power-law electron distribution with an index of 2 [85]. In accreting stellar-mass black holes, polarization fractions are typically below 10%, but can reach up to 50% during instances of optically thin emission ([68] and references therein). Thus, the upper limit on the linear polarization measured from J1023 in the high mode is consistent with the presence of an optically thick compact jet, similar to what is observed in other X-ray binaries. In contrast, the lower polarization limit during the low mode may result from partial cancellation of orthogonally polarized jet components due to misaligned magnetic fields.
Appendix G Polarimetric properties of accreting neutron stars in LMXBs
IXPE has observed several accreting NSs in LMXBs, revealing diverse polarimetric behaviors. The polarization degree of the soft X-ray emission typically remains around a few percent across all observed systems. In this section, we summarize the polarization properties of the observed systems and briefly discuss their interpretations.
G.1 Z-sources
Cyg X2.
Cyg X-2 exhibits stable spectral characteristics along the Z-track, with a notable hard X-ray excess above 30 keV during the Horizontal Branch phase linked to episodic radio jet emissions. IXPE measurements indicate a polarization degree of and a PA of , aligning with the radio jet direction. The Comptonized emission component shows a polarization degree of , also aligned with the jet, suggesting the primary polarization originates from a spreading layer on the NS surface, with potential contributions from disk reflection [48].
Sco X1.
Sco X1 is the brightest persistent extrasolar X-ray source, approaching the Eddington luminosity for a 1.4 NS. It has a 19-hour orbital period and it is located 2.13 kpc away. IXPE observations in the soft state revealed an average PD of and a PA of . The PD of the accretion disk emission was constrained to 3.2%, while that of the Comptonized component was [78].
GX 51.
GX 51 is located near the Galactic Center and is notable for radio emission from a compact jet. IXPE observations in the horizontal and normal-flaring branches of the Z-track measured PDs of and , respectively. A 20∘ variation in PA with energy was observed, likely due to differing polarization angles of the disk and Comptonization components peaking at various energies [46].
GX 3400.
GX 3400 is known for its stable spectral behavior and associated radio jet emissions. IXPE observations detected significant X-ray polarization when the source was mainly within the horizontal branch, with a PD of and a PA of in the 2–8 keV range. A variation in polarization angle with energy was measured, which was attributed to either the accretion disk and Comptonized emissions, or a mix of blackbody emission and Comptonized emission [13]. Alternatively, the low-energy polarization could be due to the accretion disk whereas the high energy polarization could arise from the corona [79].
Subsequent observations during the normal branch phase recorded a PD of and a PA of , with the PD exhibiting a marginal increase with energy. The PA is consistent with that measured in the horizontal branch. The observed polarized emission is consistent with originating from blackbody emission, Comptonized emission, or reflection features, with no significant contribution from the accretion disk [13].
G.2 Atoll sources
GS 1826238.
GS 1826238 remained in a hard spectral state until its transition to a high soft state in early 2016. IXPE observations in 2022 set a upper limit on the PD of 1.3% [25].
GX 31
GX 31 was observed with IXPE, along with NuSTAR and NICER, in August 2024. No significant X-ray polarization was detected in the 2–8 keV band, with an upper limit of 1.3% on the PD at the 99% c.l.. The X-ray spectra are well described by a model consisting of thermal emission from an accretion disk, a Comptonized emission component, and a component accounting for reflected photons off the disk. The upper limits on the polarized signal from the disk and the Comptonization component are consistent with theoretical expectations for a system with a standard disk and a Comptonizing region that likely has a geometry consistent with a spreading layer [53].
GX 99.
GX 99 exhibits a consistent bright soft state and a 4.2-hour modulation in both optical and X-ray bands. A spectro-polarimetric analysis of IXPE data detected X-ray polarization with a PD of 1.7%0.4% in the 2–8 keV band, which increased from a value consistent with null polarization in the 2–4 keV range to 3.2% in the 4–8 keV range. Spectral modeling indicated that the emission comprises thermal radiation from the accretion disk and a Comptonized component from an optically thick corona, suggesting that the polarization primarily originates from the Comptonized component [29]. A subsequent study measured a PD of 2.2%0.5% in the 4–8 keV band. Broadband spectral modeling required thermal disk emission, a Comptonized component, and a reflection component. An upper limit of 4% was found on the PD of the disk emission; both Comptonization near the NS and disk reflection were found to contribute to the observed polarization [118].
4U 162449.
4U 162449 is a persistent system known for its periodic 6–8 hour dips every 21 hours and flaring behavior above the bright persistent emission. IXPE detected an average PD of and a PA of during non-dip intervals, with PD increasing from (4–6 keV) to (6–8 keV). This suggests polarization arises from Comptonization in an extended, geometrically thin slab corona covering part of the accretion disk and disk reflection. During X-ray dips, an upper limit of 22% was established for the PD [108, 54].
G.3 Ultracompact binaries
4U 162667.
4U 162667 is a 7.66-s pulsar in an ultracompact binary with a 42-minute orbital period. IXPE observations constrained the pulse-averaged PD to 4% (95% confidence). However, phase-resolved spectro-polarimetric analysis revealed a marginal polarization detection in the power-law component with a PD of (90% confidence; [88]).
4U 1820303.
4U 1820303 is an ultra-compact low-mass X-ray binary located in the globular cluster NGC 6624 at 8 kpc. The system features a 11-minute orbital period and exhibits luminosity variations over a 170-day super-orbital cycle. Its spectrum comprises a blackbody or multicolor disk component, a Comptonization component, and an iron K line around 6.5 keV. IXPE observations measured a PD of with a PA of in the 4–7 keV band and with a PA of in the 7–8 keV band. spectro-polarimetric analysis suggests orthogonal polarization between the accretion disk and the hard spectral component from the boundary/spreading layer [40].
G.4 Transient systems
XTE J1701462.
This transient system experienced an outburst in September 2022, and was observed twice by IXPE with a 10-day interval. The initial observation was performed when the source was in the horizontal branch and recorded an average PD of , increasing from (2–4 keV) to (4–8 keV). The subsequent observation, taken in the normal branch, showed a reduced average PD of 0.6%. Spectral modeling incorporated an unpolarized thermal emission from the accretion disk and a Comptonized component with a PD of , indicating the Comptonized emission likely originates from a spreading or boundary layer above the NS surface, decreasing at later outburst stages [32, 72, 133]. A recent time-resolved polarimetric analysis revealed a stable PD of 2% in the second observation along with significant variations in the PA (yielding depolarization in the time-averaged data and hence the smaller average PD). These results point to rapid changes in the geometry of the corona, possibly from a slab geometry to a more vertically extended spreading layer geometry [134].
SRGA J144459.2604207.
SRGA J144459.2604207 is an accreting MSP detected in outburst in February 2024. Over a 10-day monitoring period, IXPE measured an average PD of in the 2–8 keV band, with a PA of . The polarization exhibited significant energy dependence, reaching between 3 and 6 keV, while remaining below 2% (90% c.l.) in the 2–3 keV range. No significant variability in the polarization properties was detected across spin and orbital phases. The observations are consistent with a moderately fan-beam emission originating from two localized hotspots with a small magnetic obliquity of . An upper limit on the PD of 8.5% (90% c.l.) was set during type I bursts [99].
G.5 Other systems
GX 131. GX 131 is a bright, persistent Galactic type I X-ray burster. IXPE observations in October 2023 detected a PD of (2–8 keV) with a PA of . A follow-up observation in February 2024 measured an overall PD of and a PA of . The polarization was attributed to a combination of Comptonization in the boundary or spreading layer and scattering from the accretion disk wind [17, 15].